OpenStudy (enjoymeicecold):

NEED HELP A cone-shaped paper cup can hold 35 cubic inches of food. If the radius of the cone is 2.4 inches, what is the height of the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 5.81 inches 6.57 inches 7.89 inches 8.89 inches

OpenStudy (acannell):

5.81 inches :)

OpenStudy (enjoymeicecold):

i need help on how u got to that answer

OpenStudy (acannell):

Oh okay one sec.

OpenStudy (acannell):

1/3 × pi × r^2 × h 35 = 1/3 × 3.14 × 2.4^2 × h 35 = 1/3 × 3.14 × 5.76 × h 35 = 6.0288h 35 / 6.0288 = ?

OpenStudy (acannell):

Do you understand?

OpenStudy (enjoymeicecold):

oh ok thx

OpenStudy (acannell):

Your welcome!! :)

OpenStudy (terryhall1999):

8.89 inches